OpenStudy (anonymous):

A 30 foot rope that is used to hold a boat near a dock forms a 32 degree with the level of the water and is connected to the boat at this level. What is the height of the dock to the nearest tenth of a foot?

OpenStudy (dumbcow):

sin(32) = h/30 h = 30*sin(32)

OpenStudy (anonymous):

Two forces of 130 and 150 pound yield a resultant force of 180 pounds. Find, to the nearest tenth of a degree, the angle between the original two forces.

OpenStudy (dumbcow):

im not 100% on this one i used law of cosines and found the angle to be 100.3 degrees

OpenStudy (anonymous):

i would have to find angle a and c and add them together correct?

OpenStudy (anonymous):

or would i be finding angle b?

OpenStudy (dumbcow):

correct between 130 and 180 i found angle 55.06 between 150 and 180 i found angle 45.27

OpenStudy (anonymous):

wait are you sure it isnt angle b?

OpenStudy (dumbcow):

whats your reference is angle c opposite of 180? if it is then yes your right add a and b

OpenStudy (anonymous):

you start off with a parallelogram correct?

OpenStudy (dumbcow):

yes with 180 vector across middle

OpenStudy (anonymous):

yes and the top triangle i labeled bottom left as A top left as B and top right as C

OpenStudy (dumbcow):

oh ok then, add A and C to find angle between 130 and 150 vectors

OpenStudy (dumbcow):

my A was 55 my C was 45